OpenStudy (anonymous):

Solve square root of 2x + 1 = x + 1 for x. a. No Solution b.±2 c.±3 d.0

OpenStudy (anonymous):

Is the square root over the 2x only? like this\[\sqrt{2x}\] or is the 1 under the squre root also?

OpenStudy (anonymous):

the one is under the square root too

OpenStudy (anonymous):

There is a trick to get rid of the radical or 'square root'; do you know how to do that?

OpenStudy (anonymous):

No not too shabby with math

OpenStudy (anonymous):

There is a law \[(\sqrt{a})^{2}=a\] if you square a square root, it becomes that thing inside the square root. You can square both sides of your equation.

OpenStudy (anonymous):

And the answer is?

OpenStudy (anonymous):

square both sides and solve it

OpenStudy (anonymous):

\[\sqrt{2x+1}=x+1\] \[(\sqrt{2x+1})^{2}= (x+1)^{2}\]